# Estimating stationary mass, frequency by frequency

### v2

@@ 4 Main results
Because $\sum_{u=1}^{4\tau-2}{\mathbb{E}}\left[M^{\pi}_{u}\right]\leq 1$, this recovers the minimax rate of $\mathcal{O}(\sqrt{\tau/n})$ given by Pananjady et al. (2024). Moreover, the bound above is strictly stronger, since the RHS can be significantly smaller when the quantities $\{\mathbb{E}[M^{\pi}_{u}]\}_{u=0}^{4\tau-2}$ are small. Note that these are the expected count probability masses for the stationary distribution $\pi$ — if the small count mass of the stationary distribution is small (i.e. most symbols appear many times), then our bound shows that WingIt incurs a smaller error for missing mass estimation. This adaptivity is a distinct advantage of our approach. In contrast, the $\mathsf{MSE}$ bound provided in Pananjady et al. (2024) captures the correct scaling in terms of $\tau$ and $n$, but is not adaptive to this specific behavior of the stationary distribution $\pi$. This adaptivity property plays a central role in our final bound in Theorem 1, since the errors incurred across all the $\zeta$ values are aggregated to produce the claimed bound on TV error.
@@ 5.2 Proof of Theorem 1
We prove Theorem 1 from Propositions 1 and 2, which in turn are proved in Sections 5.4 and 5.5. Denote $\widehat{M}^{\textsc{u}}$ as the unnormalized estimator, i.e. $\widehat{M}^{\textsc{u}}:=\widehat{M}(\tau;\overline{\zeta})\cdot\nu$, with
@@ 5.3 Proof of Lemma 2
We now construct a new auxiliary stochastic process as follows. We choose the tuple $(Y_{1},\ldots,Y_{\tau})$ independently from everything else, and according to the law of $(U_{1},\ldots,U_{\tau})$. Similarly, choose the blocks $(Y_{\tau+1},\ldots,Y_{2\tau}),\ldots,(Y_{n-\tau+1},\ldots,Y_{n})$ IID from the same distribution. In words, at the start of each interval $S_{k}$, we restart the process from a state sampled from the stationary distribution $\pi$. Let us denote this sequence as
@@
Proof of claim (25): Executing the empirical Bernstein argument for the auxiliary sequence [^4] Note that in this case, we do not need to split over odd and even subsequences; just a direct application of the empirical Bernstein bound suffices and we obtain a slightly better constant. $Y^{n}$, we have
@@ 5.4 Proof of Proposition 1
Recall that Proposition 1 bounds the $\ell_{1}$ risk between between the plug-in estimator $\widehat{M}_{\textsc{PI},\zeta}$ and and the true count probability mass $M^{\pi}_{\zeta}$ for sufficiently large values of count $\zeta$. Also recall the shorthand notation $\tau_{0}=\mathsf{t_{mix}}(\epsilon/n^{2})$. We denote by $\mathcal{X}_{\zeta}$ the set of all symbols appearing exactly $\zeta$ times in $X^{n}$ and by $\mathcal{X}_{\geq\zeta}$ the set of all symbols appearing at least $\zeta$ times in $X^{n}$. Clearly, we have $\mathcal{X}_{\zeta}\subseteq\mathcal{X}_{\geq\zeta}$. Using this fact and the triangle inequality, we have
@@ 5.5 Proof of Proposition 2
In this section we bound the error of the WingIt estimator for small frequencies. We first derive bounds on the mean squared error ($\mathsf{MSE}$) for each $\zeta$, and then use these to obtain bounds on the $\ell_{1}$ error. Throughout this proof, note that $\tau$ should be thought of as an arbitrary integer that divides $n$, with $n_{0}=n/(2\tau)$.
@@ 5.5.2 Proof of Lemma 5
First, we show an upper bound on $A_{1}$. It is convenient to introduce variables $u$ and $v$ that count the number of occurrences of $Y$ in $\mathcal{B}_{j}$ and $\mathcal{B}_{k}$ respectively. Then, we have
@@ 6 Discussion
We have provided a flexible estimator and analysis of the vector of count probabilities $M^{\pi}(X^{n})$ (and therefore, stationary distribution estimation) of any stationary $\alpha$-mixing stochastic process. An explicit construction for the IID case (Acharya et al., 2013) reveals that our estimation error rate is sharp in its dependence on $n$; we conjecture that the dependence on $\mathsf{t_{mix}}$ is tight for our estimator, but showing this remains open. While we have obtained high-probability bounds on the error of the plug-in estimator, our bounds on the WingIt are in expectation — obtaining corresponding high-probability bounds would be an interesting follow-up problem. Finally, our approach heavily uses the fact that TV decomposes as an $\ell_{1}$ norm of frequency-by-frequency errors; obtaining corresponding analysis for the KL-divergence is an intriguing direction for future work.
@@ A.3 Surrogate Process Lemmas
Let $n_{0}=n/\tau$. For each $j\in[n_{0}]$, define the stochastic process $(X^{\prime}_{k})_{k\in\mathcal{D}_{j\tau}}$ as a $|\mathcal{D}_{j\tau}|$-length sequence and initial state sampled from the distribution $\pi$ and independently of everything else. Thus the tuple $(X^{\prime}_{1},\ldots,X^{\prime}_{\tau})$ is independent of everything else, and according to the law of $(X_{1},\ldots,X_{\tau})$. Similarly, choose the blocks $(X^{\prime}_{\tau+1},\ldots,X^{\prime}_{2\tau}),\cdots,(X^{\prime}_{n-\tau+1},\ldots,X^{\prime}_{n})$ i.i.d. from the same distribution. We also assume the stationary stochastic process $X_{t}$ is exponentially $\alpha$-mixing with parameters $\mu$ and $\rho$ (see Eq. (1b)). Now, define the stochastic processes
@@
$\widetilde{W}_{j}^{(k)}=\left(\bigoplus_{k^{\prime}=1}^{k}X_{\mathcal{B}_{\ell_{k^{\prime}}}}\right)\bigoplus\left(\bigoplus_{k^{\prime}=k+1}^{s}X^{\prime}_{\mathcal{B}_{\ell_{k^{\prime}}}}\right).$
@@
$\leq\sup_{S_{1},\ldots,S_{s}\subseteq\mathcal{X}}|\mathbb{P}(X_{\mathcal{B}_{\ell_{1}}}\in S_{1},\ldots,X_{\mathcal{B}_{\ell_{k}}}\in S_{k})\mathbb{P}(X_{\mathcal{B}_{\ell_{k+1}}}\in S_{k+1})$
$\qquad\qquad\qquad-\mathbb{P}(X_{\mathcal{B}_{\ell_{1}}}\in S_{1},\ldots,X_{\mathcal{B}_{\ell_{k}}}\in S_{k},X_{\mathcal{B}_{\ell_{k+1}}}\in S_{k+1})|$

### v3

@@ 4 Main results
Because $\sum_{u=1}^{4\tau-2}{\mathbb{E}}\left[M^{\pi}_{u}\right]\leq 1$, this recovers the minimax rate of $\mathcal{O}(\sqrt{\tau/n})$ given by Chandra et al. (2022). Moreover, the bound above is strictly stronger, since the RHS can be significantly smaller when the quantities $\{\mathbb{E}[M^{\pi}_{u}]\}_{u=0}^{4\tau-2}$ are small. Note that these are the expected count probability masses for the stationary distribution $\pi$ — if the small count mass of the stationary distribution is small (i.e. most symbols appear many times), then our bound shows that WingIt incurs a smaller error for missing mass estimation. This adaptivity is a distinct advantage of our approach. In contrast, the $\mathsf{MSE}$ bound provided in Pananjady et al. (2024) captures the correct scaling in terms of $\tau$ and $n$, but is not adaptive to this specific behavior of the stationary distribution $\pi$. This adaptivity property plays a central role in our final bound in Theorem 1, since the errors incurred across all the $\zeta$ values are aggregated to produce the claimed bound on TV error.
5 Numerical experiments
In this section, we provide a set of simulations on synthetically constructed Markov chains in order to demonstrate the effectiveness of the proposed estimator. We first describe the experimental set-up and discuss some of the estimators used in practice. We then compare the performance of these estimators against the proposed hybrid estimator.
5.1 Experimental Setup
We perform experiments on samples that we generate from different types of $\alpha$-mixing processes. Specifically, we instantiate the stationary distribution of the mixing sequence from one of five different distributions. These distributions include: the uniform distribution, a step distribution with half of the symbols occurring with probability $1/(2|\mathcal{X}|)$ and the other half of symbols occurring with probability $3/(2|\mathcal{X}|)$, a Dirichlet distribution with the concentration parameter $1.5$, and Zipf distributions with parameters $1.1$ and $1.5$. The Zipf distribution with parameter $a$ assigns the probability $p(i)\propto i^{-a}$ for all $i\in\mathbb{N}$. We then layer a random duplication model on the i.i.d. samples generated from the stationary distribution to obtain the final $\alpha$-mixing sequence. This process is described below. For notational convenience, we define $\mathsf{T_{mix}}\coloneqq\mathsf{t_{mix}}(1/4)$, where $\mathsf{t_{mix}}(\epsilon)$ is defined in Eq. (2).
  1. 1.
    First, we generate $n_{0}$ i.i.d. samples from the underlying stationary distribution. The number of symbols is set as $|\mathcal{X}|=n_{0}$. Thus, as we increase the number of samples, the number of symbols increases proportionally.
  2. 2.
    We then use a geometric duplication model in order to generate a mixing sequence from the i.i.d. samples. Specifically, for each sample in the i.i.d. sequence, we draw a geometric random variable with mean $n_{0}^{c}$ (for some $c>0$), which determines how many times that symbol is duplicated in the final sequence. This processcan be viewed as a draw from a sticky Markov chain (Chandra et al., 2022) in which the probability of observing an i.i.d. sample from the stationary distribution is given by $1/n_{0}^{c}$. The mixing time, $\mathsf{T_{mix}}$, for this sequence is given by $\mathsf{T_{mix}}=\mathcal{O}(n_{0}^{c})$ (Pananjady et al., 2024). Since the duplication factors are random, the total length of the resulting sequence, denoted as $n$, is also a random variable with expected value $\mathbb{E}[n]=n_{0}^{c+1}$.
We compare the performance of our proposed estimator with several existing estimators that were originally designed for i.i.d. sequences. Specifically, we include the Good–Turing + Plug-in estimator analyzed by Acharya et al. (2013), as well as a family of popular add-$\beta$ estimators defined as:
$\widehat{M}_{\zeta}(\beta)=\frac{\varphi_{\zeta}(\zeta+\beta(\zeta))}{\nu},$
where $\nu$ is a normalizing constant chosen to ensure that the estimated probabilities sum to $1$. Different choices of the function $\beta(\zeta)$ yield well-known estimators:
  • Laplace estimator (Laplace, 1814): $\beta(\zeta)=1$ for all $\zeta$,
  • Krichevsky–Trofimov estimator (Krichevsky and Trofimov, 1981): $\beta(\zeta)=0.5$ for all $\zeta$,
  • Braess–Sauer estimator (Braess and Sauer, 2004): $\beta(0)=0.5$, $\beta(1)=1$, and $\beta(\zeta)=0.75$ for all $\zeta>1$.
An important detail is that the normalization constant $\nu$ for the add-$\beta$ estimators depends on the alphabet size $|\mathcal{X}|$. For the uniform, step, and Dirichlet distributions, the alphabet size is finite and set to $|\mathcal{X}|=n_{0}$. However, for the Zipf distribution, the alphabet is the set of all natural numbers, so that $|\mathcal{X}|=\infty$. To address this in our simulations, we approximate the alphabet size for the Zipf distribution using the number of unique symbols observed in the samples. This approximation is reasonable because the missing mass—corresponding to mass placed on unobserved symbols—is small with high probability for Zipf distribution and has a negligible effect on the TV distance.
We compare estimator performance by plotting the TV distance between the true and estimated vectors of count probabilities for each method. We present simulation results for three different values of $c$, the parameter controlling the mixing time of the sequences. All plots are averaged over $50$ independent instances, and the shaded regions around each curve represent the standard deviation across these runs.
(a) Uniform
(b) Dirichlet
(c) Zipf with parameter $1.1$
(d) Zipf with parameter $1.5$
(e) Step
Figure 1: $c=0.0$ corresponding to the IID sequence for $n_{0}$ samples, averaged over $50$ instances.
(a) Uniform
(b) Dirichlet
(c) Zipf with parameter $1.1$
(d) Zipf with parameter $1.5$
(e) Step
Figure 2: $c=0.2$ corresponding to a mixing sequence with mixing time $\mathsf{T_{mix}}=\mathcal{O}(n_{0}^{0.2})$ and expected number of samples ${\mathbb{E}}[n]=n_{0}^{1.2}$, averaged over $50$ instances.
(a) Uniform
(b) Dirichlet
(c) Zipf with parameter $1.1$
(d) Zipf with parameter $1.5$
(e) Step
Figure 3: $c=0.2$ corresponding to a mixing sequence with mixing time $\mathsf{T_{mix}}=\mathcal{O}(n_{0}^{0.33})$ expected number of samples ${\mathbb{E}}[n]=n_{0}^{1.33}$, averaged over $50$ instances.
5.2 Results
First, in Figure 1 we plot the results for the case $c=0.0$. Since the duplication factor for each sample in this case is simply $1$, this represents the case of executing the estimators on an i.i.d. sequence and is a convenient reference point for evaluation. In this case, the WingIt+Plug-in estimator is exactly equal to the Good–Turing+Plug-in estimator and thus the performance of these estimators overlap; we only include the evaluation of the WingIt+Plug-in estimator to avoid redundancy. This estimator performs well on all distributions and is significantly better than the add-$\beta$ estimators on all base distributions.
Figure 2 shows the results for $c=0.2$, where the data is no longer i.i.d. The window size for the WingIt+Plug-in estimator is set to $\tau=10\times n_{0}^{c}$, which is on the order of the mixing time of the sequence. In this setting, the WingIt+Plug-in estimator continues to outperform all competing methods across all stationary distributions considered. It exhibits consistent estimation behavior, with the TV distance from the true count probabilities decreasing steadily as the number of samples increase. In contrast, the add-$\beta$ estimators fail to achieve consistent estimation across the different base distributions.
Finally, Figure 3 presents the results for $c=0.33$, where the data is even more dependent. Once again, we observe that the WingIt+Plug-in estimator significantly outperforms the other estimators, highlighting its effectiveness in dealing with temporal dependencies in the data.
@@ 6.2 Proof of Theorem 1
We prove Theorem 1 from Propositions 1 and 2, which in turn are proved in Sections 6.4 and 6.5. Denote $\widehat{M}^{\textsc{u}}$ as the unnormalized estimator, i.e. $\widehat{M}^{\textsc{u}}:=\widehat{M}(\tau;\overline{\zeta})\cdot\nu$, with
@@ 6.3 Proof of Lemma 2
We now construct a new auxiliary stochastic process as follows. We choose the tuple $(Y_{1},\ldots,Y_{\tau})$ independently of everything else, and according to the law of $(U_{1},\ldots,U_{\tau})$. Similarly, choose the blocks $(Y_{\tau+1},\ldots,Y_{2\tau}),\ldots,(Y_{n-\tau+1},\ldots,Y_{n})$ IID from the same distribution. In words, at the start of each interval $S_{k}$, we restart the process from a state sampled from the stationary distribution $\pi$. Let us denote this sequence as
@@
Proof of claim (25): Executing the empirical Bernstein argument for the auxiliary sequence [^4] Note that in this case, we do not need to split over odd and even subsequences; just a direct application of the empirical Bernstein bound suffices, and we obtain a slightly better constant. $Y^{n}$ we have
@@ 6.4 Proof of Proposition 1
Recall that Proposition 1 bounds the $\ell_{1}$ risk between the plug-in estimator $\widehat{M}_{\textsc{PI},\zeta}$ and the true count probability mass $M^{\pi}_{\zeta}$ for sufficiently large values of count $\zeta$. Also recall the shorthand notation $\tau_{0}=\mathsf{t_{mix}}(\epsilon/n^{2})$. We denote by $\mathcal{X}_{\zeta}$ the set of all symbols appearing exactly $\zeta$ times in $X^{n}$ and by $\mathcal{X}_{\geq\zeta}$ the set of all symbols appearing at least $\zeta$ times in $X^{n}$. Clearly, we have $\mathcal{X}_{\zeta}\subseteq\mathcal{X}_{\geq\zeta}$. Using this fact and the triangle inequality, we have
@@ 6.5 Proof of Proposition 2
In this section, we bound the error of the WingIt estimator for small frequencies. We first derive bounds on the mean squared error ($\mathsf{MSE}$) for each $\zeta$, and then use these to obtain bounds on the $\ell_{1}$ error. Throughout this proof, note that $\tau$ should be thought of as an arbitrary integer that divides $n$, with $n_{0}=n/(2\tau)$.
@@ 6.5.2 Proof of Lemma 5
First, we show an upper bound on $A_{1}$. It is convenient to introduce variables $u$ and $v$ that count the number of occurrences of $Y$ in $\mathcal{B}_{j}$ and, $\mathcal{B}_{k}$ respectively. Then, we have
@@ 7 Discussion
We have provided a flexible estimator and analysis of the vector of count probabilities $M^{\pi}(X^{n})$ (and therefore, stationary distribution estimation) of any stationary exponentially $\alpha$-mixing stochastic process. An explicit construction for the IID case (Acharya et al., 2013) reveals that our estimation error rate is sharp in its dependence on $n$; we conjecture that the dependence on $\mathsf{t_{mix}}$ is tight for our estimator, but showing this remains open. While we have obtained high-probability bounds on the error of the plug-in estimator, our bounds on the WingIt are in expectation — obtaining corresponding high-probability bounds would be an interesting follow-up problem. Finally, our approach heavily uses the fact that TV decomposes as an $\ell_{1}$ norm of frequency-by-frequency errors; obtaining corresponding analysis for the KL-divergence is an intriguing direction for future work.
@@ A.3 Surrogate Process Lemmas
Let $n_{0}=n/\tau$. For each $j\in[n_{0}]$, define the stochastic process $(X^{\prime}_{k})_{k\in\mathcal{D}_{j\tau}}$ as a $|\mathcal{D}_{j\tau}|$-length sequence and initial state sampled from the distribution $\pi$ and independently of everything else. Thus, the tuple $(X^{\prime}_{1},\ldots,X^{\prime}_{\tau})$ is independent of everything else, and according to the law of $(X_{1},\ldots,X_{\tau})$. Similarly, choose the blocks $(X^{\prime}_{\tau+1},\ldots,X^{\prime}_{2\tau}),\cdots,(X^{\prime}_{n-\tau+1},\ldots,X^{\prime}_{n})$ i.i.d. from the same distribution. We also assume the stationary stochastic process $X_{t}$ is exponentially $\alpha$-mixing with parameters $\mu$ and $\rho$ (see Eq. (1b)). Now, define the stochastic processes
@@
$\widetilde{W}_{j}^{(k)}=\Big{(}\bigoplus_{k^{\prime}=1}^{k}X_{\mathcal{B}_{\ell_{k^{\prime}}}}\Big{)}\bigoplus\Big{(}\bigoplus_{k^{\prime}=k+1}^{s}X^{\prime}_{\mathcal{B}_{\ell_{k^{\prime}}}}\Big{)}.$
@@
$\leq\sup_{S_{1},\ldots,S_{s}\subseteq\mathcal{X}}|\mathbb{P}(X_{\mathcal{B}_{\ell_{1}}}\in S_{1},\ldots,X_{\mathcal{B}_{\ell_{k}}}\in S_{k})\mathbb{P}(X_{\mathcal{B}_{\ell_{k+1}}}\in S_{k+1})-\mathbb{P}(X_{\mathcal{B}_{\ell_{1}}}\in S_{1},\ldots,X_{\mathcal{B}_{\ell_{k}}}\in S_{k},X_{\mathcal{B}_{\ell_{k+1}}}\in S_{k+1})|$